# A Systematic Evaluation of Self-Supervised Learning for Label-Efficient Sleep Staging with Wearable EEG

### v1

@@ 2.3 Sleep Scoring with Wearable EEG Devices
Recent research has increasingly focused on evaluating wearable EEG devices for sleep staging, primarily by benchmarking their performance against the gold standard of PSG. A systematic review by De Gans et al. [10] analyzed 60 studies encompassing 34 unique EEG-based wearables, highlighting a surge in interest, with over half of the reviewed publications appearing from 2020 onward. The reviewed devices demonstrated promising accuracy in sleep staging, often approaching that of PSG. However, the vast majority of studies were conducted in controlled, clinical settings, with only 12% of them validating wearables in home environments. Moreover, all approaches relied on fully labeled datasets, leaving the potential of label-efficient training paradigms, such as SSL, largely unexplored. Our work addresses this gap by systematically evaluating SSL for sleep staging using wearable EEG data.
@@ 6.3 Applications and Future Directions
In this paper, we address the challenge posed by the large volumes of unlabeled EEG data generated through the widespread adoption of wearable devices in the context of automatic sleep staging. To leverage this unlabeled data, we evaluate SSL techniques as a pre-training step within conventional supervised learning pipelines. From a practical standpoint, SSL can be incorporated into existing sequence-to-sequence sleep staging frameworks with minimal overhead, requiring no changes to the model architecture or downstream task. As a result, SSL consistently improves classification performance over purely supervised baselines, and enhances generalization and cross-dataset transferability. In scenario 1, SSL pre-training with only 7.5% of labeled data achieves comparable accuracy to supervised learning with 15% of labels, representing twice the labeling effort, while using 20% of labels with SSL results in only a -2% difference compared to fully supervised training with 100% of the dataset. This makes the development of sleep scoring systems more label-efficient and cost-effective, reducing dependence on manual annotations, which are expensive to obtain and subject to inter-scorer variability. Our findings using wearable EEG data align with prior evaluations conducted on PSG data [20], particularly in terms of performance under varying data availability and domain generalization, further reinforcing the viability of SSL for wearable sleep staging.
Based on the evidence provided, the application of SSL in sleep scoring pipelines proves largely valuable when labeled data is scarce, a substantial volume of unlabeled EEG sleep data is available, and baseline performance falls below the lower bound of the inter-scorer agreement (< 80%), which we identify as the medical-grade benchmark. Beyond this, SSL offers valuable opportunities in commercial consumer sleep technologies, enabling regular model updates without requiring manual scoring. Its strong generalization also makes it well-suited for distributed and federated learning setups. Additionally, SSL can support continual learning and personalized model adaptation using limited user-specific data. Importantly, the SSL techniques employed are agnostic to the domain, making them applicable to other clinical areas where labeled data is even harder to obtain.

### v2

@@ 2.3 Sleep Scoring with Wearable EEG Devices
Recent research has increasingly focused on evaluating wearable EEG devices for sleep staging, primarily by benchmarking their performance against the gold standard of PSG. A systematic review by De Gans et al. [10] analyzed 60 studies encompassing 34 unique EEG-based wearables, highlighting a surge in interest, with over half of the reviewed publications appearing from 2020 onward. The reviewed devices demonstrated promising accuracy in sleep staging, often approaching that of PSG. However, the vast majority of studies were conducted in controlled, clinical settings, with only 12% of them validating wearables in home environments. Moreover, most approaches relied on fully labeled datasets, leaving the potential of label-efficient training paradigms, such as SSL, largely unexplored. Khalique et al. [50] provided an important first step toward integrating SSL with portable EEG for in-home sleep scoring. Nonetheless, the scope of their work is limited to a single SSL method and training setting. Our study conducts the first comprehensive evaluation of SSL for wearable EEG sleep staging, comparing multiple state-of-the-art SSL methods across different evaluation scenarios using both clinically annotated and large-scale unlabeled home recordings to analyze label efficiency, representation quality, and cross-dataset generalization.
@@ 6.3 Applications and Future Directions
In this paper, we address the challenge posed by the large volumes of unlabeled EEG data generated through the widespread adoption of wearable devices in the context of automatic sleep staging. To leverage this unlabeled data, we evaluate SSL techniques as a pre-training step within conventional supervised learning pipelines. From a practical standpoint, SSL can be incorporated into existing sequence-to-sequence sleep staging frameworks with minimal overhead, requiring no changes to the model architecture or downstream task. As a result, SSL consistently improves classification performance over purely supervised baselines, and enhances generalization and cross-dataset transferability. In scenario 1, SSL pre-training with only 7.5% of labeled data achieves comparable accuracy to supervised learning with 15% of labels, representing twice the labeling effort, while using 20% of labels with SSL results in only a -2% difference compared to fully supervised training with 100% of the dataset. This makes the development of sleep scoring systems more label-efficient and cost-effective, reducing dependence on manual annotations, which are expensive to obtain and subject to inter-scorer variability. Our findings using wearable EEG data align with prior evaluations conducted on PSG data [20], particularly in terms of performance under varying data availability and domain generalization, further reinforcing the viability of SSL for wearable sleep staging. Moreover, our work serves as a comprehensive benchmark for wearable EEG sleep scoring and extends beyond prior studies [50] by providing the first systematic evaluation of SSL in this domain. We evaluate multiple SSL techniques across diverse training scenarios to rigorously assess label efficiency, representation quality, and cross-dataset generalization, establishing a foundational reference for future research on SSL with wearable EEG.
Based on the evidence provided, the application of SSL in sleep staging pipelines proves largely valuable when labeled data is scarce, a substantial volume of unlabeled EEG sleep data is available, and baseline performance falls below the lower bound of the inter-scorer agreement (< 80%), which we identify as the medical-grade benchmark. Beyond this, SSL offers valuable opportunities in commercial consumer sleep technologies, enabling regular model updates without requiring manual scoring. Its strong generalization also makes it well-suited for distributed and federated learning setups. Additionally, SSL can support continual learning and personalized model adaptation using limited user-specific data. Importantly, the SSL techniques employed are agnostic to the domain, making them applicable to other clinical areas where labeled data is even harder to obtain.